OpenStudy (anonymous):

test in an hour! please help...sally transfers 34 credit hours with a gpa of 4.0 from ST university to BT university where she completes another 12 hrs with a gpa of 2.4. What is her final cumulative gpa rounded to 3 decimal places?

OpenStudy (anonymous):

These kind of questions you eliminate words, other stuff, zoom in on what question asks. Question as about gpa. What are the two gpa?

OpenStudy (anonymous):

4.0 and 2.4

OpenStudy (anonymous):

What is final cumulative gpa?

OpenStudy (anonymous):

i dont know:( how would you solve it?

OpenStudy (anonymous):

If I score 20 points in game 1 and 10 points in game 2. 20+10=30 30/2 games=15.000 (rounded off to 3 dec points)

OpenStudy (anonymous):

ok so its just like doing an average?

OpenStudy (anonymous):

Yes cumulative is just a fancy word that makes you sound smart

OpenStudy (anonymous):

hahaha! so it would be 4.0+2.4 = 6.4 then 6.4/2?

OpenStudy (anonymous):

Rounded to 3 decimal points

OpenStudy (anonymous):

thank you so much for your help:)

OpenStudy (anonymous):

You're going to ace this test!

OpenStudy (anonymous):

lol i hope:/